# Nested subspace learning with flags

### v1

@@ 3.2 The flag trick
Let $p\geq 2$, $0<q<p$ and $q_{1:d}:=(q_{1},q_{2},\cdots,q_{d})$ be a sequence of increasing dimensions such that $0<q_{1}<q_{2}<\cdots<q_{d}<p$. The flag trick consists in replacing a subspace learning problem of the form:
$\underset{\mathcal{S}\in\operatorname{Gr}(p,q)}{\operatorname{argmin}}\,f(\Pi_{\mathcal{S}})$ (10)
@@
$\underset{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}{\operatorname{argmin}}\,f\left(\frac{1}{d}\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}}\right).$ (11)
Except for the very particular case of PCA (Theorem 4) where $f_{X}(\Pi)=\left\lVert X-\Pi X\right\rVert_{F}^{2}$, we cannot expect to have an analytic solution to the flag problem (11); indeed, in general, subspace problems do not even have a closed-form solution as we shall see in section 4. This justifies the introduction of optimization algorithms on flag manifolds like Algorithm 1.
@@
The original idea of accumulated unexplained variance (Pennec, 2018) (and its subsequent application to several variants of PCA under the name of “flagification” (Mankovich et al., 2024)) consists in summing the subspace criteria at different dimensions, while the flag trick directly averages the orthogonal projection matrices that appear inside the objective function. While both ideas are equally worth experimenting with, we believe that the flag trick has a much wider reach. Indeed, from a technical viewpoint, the flag trick appears at the covariance level and directly yields a hierarchical reweighting of the principal subspaces. This reweighting is only indirect with the AUV—due to the linearity of the trace operator—and is not expected to happen for other methods than PCA. Notably, as we shall see in section 4, the flag trick enables to easily develop extensions of well-known methods involving PCA, like IRLS (Lerman and Maunu, 2018a) or Newton-Lanczos methods for trace ratio optimization (Ngo et al., 2012), and is closer in spirit to the statistical formulations of PCA (Szwagier and Pennec, 2024).
@@ 3.3 Multilevel machine learning
Subspace learning is often used as a preprocessing task before running a machine learning algorithm, notably to overcome the curse of dimensionality. One usually projects the data onto the optimal subspace $\mathcal{S}^{*}\in\operatorname{Gr}(p,q)$ and use the resulting lower-dimensional dataset as an input to a machine learning task like clustering, classification or regression (Bouveyron et al., 2019). Since the flag trick problem (11) does not output one subspace but a hierarchical sequence of nested subspaces, it is legitimate to wonder what to do with such a multilevel representation. In this subsection, we propose a general ensemble learning method to aggregate the hierarchical information coming from the flag of subspaces.
Let us consider a dataset $X:=[x_{1}|\dots|x_{n}]\in\mathbb{R}^{p\times n}$ (possibly with some associated labels $Y:=[y_{1}|\dots|y_{n}]\in\mathbb{R}^{m\times n}$). In machine learning, one often fits a model to the dataset by optimizing an objective function of the form $R_{X,Y}(g)=\frac{1}{n}\sum_{i=1}^{n}L(g(x_{i}),y_{i})$. With the flag trick, we get a filtration of projected data points $Z_{k}=\Pi_{\mathcal{S}_{k}^{*}}X,k\in[1,d]$ that can be given as an input to different machine learning algorithms. This yields optimal predictors $g_{k}^{*}=\operatorname{argmin}\,R_{Z_{k},Y}$ which can be aggregated via ensembling methods. For instance, voting methods choose the model with the highest performance on holdout data; this corresponds to selecting the optimal dimension $q^{*}\in q_{1:d}$ a posteriori, based on the machine learning objective. A more nuanced idea is the one of soft voting, which makes a weighted averaging of the predictions. The weights can be uniform, proportional to the performances of the individual models, or learned to maximize the performance of the weighted prediction (Perrone and Cooper, 1992). Soft voting gives different weights to the nested subspaces depending on their contribution to the ensembled prediction and therefore provides a soft measure of the relative importance of the different dimensions. In that sense, it goes beyond the classical manifold assumption stating that data has one intrinsic dimension, and instead proposes a soft blend between dimensions that is adapted to the learning objective. This sheds light on the celebrated paper of Minka for the automatic choice of dimensionality in PCA (Minka, 2000, Section 5). Many other ensembling methods are possible like gradient boosting, Bayesian model averaging and stacking. The whole methodology is summarized in Algorithm 2.
@@
$X:=[x_{1}|\dots|x_{n}]\in\mathbb{R}^{p\times n}$ a data matrix; $q_{1:d}:=(q_{1},\dots,q_{d})$ a flag signature; $f$ a subspace learning objective; (opt.) $Y:=[y_{1}|\dots|y_{n}]\in\mathbb{R}^{m\times n}$ a label matrix
${\mathcal{S}}_{1:d}^{*}\leftarrow\operatorname{argmin}_{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}\,f(\Pi_{\mathcal{S}_{1:d}})$triangleright flag trick (11) + optimization (Alg. 1)
@@
     $g_{k}^{*}\leftarrow\operatorname{fit}(\Pi_{\mathcal{S}_{k}^{*}}X,Y)$triangleright learning on $q_{k}$-dimensional projected data
     $Y_{k}^{*}\leftarrow g_{k}^{*}(\Pi_{\mathcal{S}_{k}^{*}}X)$triangleright prediction on $q_{k}$-dimensional projected data
@@
$Y^{*}\leftarrow\operatorname{ensembling}({Y}_{1}^{*},\dots,{Y}_{d}^{*})$triangleright weighted predictions
@@
Algorithm 2 is a general proposition of multilevel machine learning with flags, but many other uses of the optimal flag $\mathcal{S}_{1:d}^{*}$ are possible, depending on the application. For instance, one may directly use the reweighted data matrix $\Pi_{\mathcal{S}_{1:d}^{*}}X$ as an input to the machine learning algorithm. This enables to fit only one model instead of $d$. One can also simply analyze the projected data qualitatively via scatter plots or reconstruction plots as evoked in section 1. The nestedness will automatically bring consistency contrarily to non-nested subspace methods, and therefore improve interpretability. Finally, many other ideas can be borrowed from the literature on subspace clustering and flag manifolds (Draper et al., 2014; Launay et al., 2021; Ma et al., 2021; Mankovich et al., 2022; Mankovich and Birdal, 2023; Mankovich et al., 2024), for instance the computation of distances between flags coming from different datasets as a multilevel measure of similarity between datasets.
@@ 4.1 Outline and experimental setting
The general methodology to compare Grassmann-based methods to flag-based methods is the following one. For each experiment, we first choose a flag signature $\leavevmode\nobreak\ {q_{1\rightarrow d}:=(q_{1},\dots,q_{d})}$, then we run independent optimization algorithms on $\operatorname{Gr}(p,q_{1}),\dots,\operatorname{Gr}(p,q_{d})$ (10) and finally we compare the optimal subspaces $\mathcal{S}_{k}^{*}\in\operatorname{Gr}(p,q_{k})$ to the optimal flag of subspaces $\mathcal{S}_{1:d}^{*}\in\operatorname{Fl}(p,q_{1\rightarrow d})$ obtained via the flag trick (11). To show the nestedness issue in Grassmann-based methods, we compute the subspace distances $\Theta(\mathcal{S}_{k}^{*},\mathcal{S}_{k+1}^{*})_{k=1\dots d-1}$, where $\Theta$ is the generalized Grassmann distance of Ye and Lim (2016, Eq. (14)). It consists in the $\ell_{2}$ norm of the principal angles, which can be obtained from the singular value decomposition (SVD) of the inner-products matrices ${U_{k}}^{\top}{U_{k+1}}$, where $U_{k}\in\operatorname{St}(p,q_{k})$ is an orthonormal basis of $\mathcal{S}_{k}^{*}$.
Regarding the implementation of the steepest descent algorithm on flag manifolds (Algorithm 1), we develop a new class of manifolds in PyManOpt (Boumal et al., 2014; Townsend et al., 2016), and run their SteepestDescent algorithm. Our implementation of the Flag class is based on the Stiefel representation of flag manifolds, detailed in section 2, with the retraction being the polar retraction. For the computation of the gradient, we use automatic differentiation with the autograd package. We could derive the gradients by hand from the expressions we get, but we use automatic differentiation as strongly suggested in PyManOpt’s documentation. Finally, the real datasets and the machine learning methods used in the experiments can be found in scikit-learn (Pedregosa et al., 2011).
@@ 4.2 The flag trick for robust subspace recovery problems
Let us consider a dataset that is a union of inliers and outliers—the inliers are assumed to lie near a low-dimensional subspace $\mathcal{S}$ while the outliers live in the ambient space. The aim of robust subspace recovery (RSR) is to recover $\mathcal{S}$ (Lerman and Maunu, 2018b). In that sense, RSR is an outlier-robust extension of classical dimension reduction methods like PCA. Without further specifications, the RSR problem might not be well-posed. Therefore, the works in this domain often have to make some assumptions on the inlier and outlier distributions in order to obtain some convergence and recovery guarantees. For instance, in Lerman and Maunu (2018b), it is assumed that the inliers “fill” the lower-dimensional subspace and that the outliers are not much “aligned”; this is rigorously defined in Lerman et al. (2015) and Maunu et al. (2019) through permeance and alignment statistics. A typical generative model following those assumptions is the Haystack model introduced in Lerman et al. (2015). The Haystack model assumes an isotropic Gaussian distribution on the subspace for the inliers and an isotropic Gaussian distribution on the (full) ambient space for the outliers. A more realistic model—the generalized Haystack model—is introduced in Maunu et al. (2019) to circumvent the simplistic nature of the Haystack model. This one assumes general (anisotropic) Gaussian distributions for the inliers and outliers. This makes the learning harder, since the anisotropy may keep the inliers from properly permeating the low-dimensional subspace—as discussed in section 1. Therefore, one has to make some stronger assumptions on the inlier-outlier ratio and the covariance eigenvalues distributions to derive some convergence and recovery guarantees.
@@
The inlier distribution in the Haystack model follows the isotropic PPCA model (Bouveyron et al., 2007, 2011), while it follows the PPCA model (Tipping and Bishop, 1999) in the case of the generalized Haystack model. Both models are a special case of the principal subspace analysis models (Szwagier and Pennec, 2024). However, as argued in Szwagier and Pennec (2024), while the Haystack model is parameterized with Grassmannians, the generalized Haystack model—which has more degrees of freedom accounting for the anisotropy—is parameterized with Stiefel manifolds. Therefore, from a statistical modeling perspective, it only makes sense to conduct subspace learning experiments on the Haystack model and not the generalized one.

### v2

@@ 3.2 The Flag Trick
Let $p\geq 2$, $0<q<p$ and $q_{1:d}\coloneqq(q_{1},q_{2},\dots,q_{d})$ be a sequence of increasing dimensions such that $0<q_{1}<q_{2}<\dots<q_{d}<p$. The flag trick consists in replacing a subspace learning problem of the form:
$\operatorname*{arg\,min}_{\mathcal{S}\in\operatorname{Gr}(p,q)}f(\Pi_{\mathcal{S}})$ (10)
@@
$\operatorname*{arg\,min}_{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}\,f\left(\Pi_{\mathcal{S}_{1:d}}\right),$ (11)
where $\penalty 10000\ {\Pi_{\mathcal{S}_{1:d}}\coloneqq\frac{1}{d}\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}}}$ is the average multilevel projector.
Let us note that for d=1d=1—i.e., with signatures of the form $q_{1:d}=(q)$—the flag-tricked problem (11) is equivalent to the original subspace problem (10). For $d>1$, the flag trick yields new objective criteria that we cannot a priori relate to the original subspace learning problem. Except for the very particular case of PCA where $f_{X}(\Pi)=\left\lVert X-\Pi X\right\rVert_{F}^{2}$ (Theorem 8), we cannot expect to have an analytic solution to the flag problem (11); indeed, in general, subspace problems do not even have a closed-form solution as we shall see in Section 4. This justifies the introduction of optimization algorithms on flag manifolds like Algorithm 1.
@@
The original idea of accumulated unexplained variance (pennec_barycentric_2018) (and its subsequent application to several variants of PCA under the name of “flagification” (mankovich_fun_2024)) consists in summing the subspace criteria at different dimensions, while the flag trick directly averages the orthogonal projection matrices that appear inside the objective function. While both ideas are equally worth experimenting with, we believe that the flag trick has a much wider practical reach. Indeed, from a technical viewpoint, the flag trick appears at the covariance level and directly yields a hierarchical reweighting of the principal subspaces (see end of Section 3.1). This reweighting is only indirect with the AUV—due to the linearity of the trace operator—and is not expected to occur beyond PCA (flagification). Notably, as we shall see in Section 4, the flag trick enables to easily develop extensions of well-known methods involving PCA, like IRLS (lerman_fast_2018) or Newton-Lanczos methods for trace ratio optimization (ngo_trace_2012), it less likely reaches singularities of the objective function (see, notably, end of Section 4.2.1) and it is closer in spirit to the statistical formulations of PCA (szwagier_curse_2025).
@@ 3.3 Multilevel Machine Learning
Subspace learning is often used as a preprocessing task before running a machine learning algorithm, notably to overcome the curse of dimensionality. One usually projects the data onto the optimal subspace $\mathcal{S}^{*}\in\operatorname{Gr}(p,q)$ and use the resulting lower-dimensional dataset as an input to a machine learning task like clustering, classification or regression (bouveyron_model-based_2019). Since the flag trick problem (11) does not output one subspace but a hierarchical sequence of nested subspaces, it is legitimate to wonder what to do with such a multilevel representation. In this subsection, we propose a general ensemble learning method to aggregate the hierarchical information coming from the flag of subspaces.
Let us consider a dataset $X\coloneqq\begin{bmatrix}x_{1}&\cdots&x_{n}\end{bmatrix}\in\mathbb{R}^{p\times n}$ (possibly with some associated labels $Y\coloneqq\begin{bmatrix}y_{1}&\cdots&y_{n}\end{bmatrix}\in\mathbb{R}^{m\times n}$). In machine learning, one often fits a model to the dataset by optimizing an objective function of the form $R_{X,Y}(g)=\frac{1}{n}\sum_{i=1}^{n}L(g(x_{i}),y_{i})$. With the flag trick, we get a filtration of projected data points $Z_{k}=\Pi_{\mathcal{S}_{k}^{*}}X,k\in\{1,\dots,d\}$ that can be given as an input to different machine learning algorithms. This yields optimal predictors $g_{k}^{*}=\operatorname{argmin}_{g}\,R_{Z_{k},Y}(g)$ which can be aggregated via ensembling methods. For instance, voting methods choose the model with the highest performance on holdout data; this corresponds to selecting the optimal dimension $q^{*}\in q_{1:d}$ a posteriori, based on the machine learning objective. A more nuanced idea is the one of soft voting, which makes a weighted averaging of the predictions. The weights can be uniform, proportional to the performances of the individual models, or learned to maximize the performance of the weighted prediction (perrone_when_1992). Soft voting gives different weights to the nested subspaces depending on their contribution to the ensembled prediction and therefore provides a soft measure of the relative importance of the different dimensions. In that sense, it goes beyond the classical manifold assumption stating that data has one intrinsic dimension, and instead proposes a soft blend between dimensions that is adapted to the learning objective. This sheds light on the celebrated paper of Minka for the automatic choice of dimensionality in PCA (minka_automatic_2000, Section 5). Many other ensembling methods are possible like gradient boosting, Bayesian model averaging and stacking. The whole methodology is summarized in Algorithm 2 and illustrated in Figure 2.
@@
$X\coloneqq\begin{bmatrix}x_{1}&\cdots&x_{n}\end{bmatrix}\in\mathbb{R}^{p\times n}$ a data matrix; $q_{1:d}\coloneqq(q_{1},\dots,q_{d})$ a flag signature; $f$ a subspace learning objective; (opt.) $Y\coloneqq\begin{bmatrix}y_{1}&\cdots&y_{n}\end{bmatrix}\in\mathbb{R}^{m\times n}$ a label matrix
${\mathcal{S}}_{1:d}^{*}\leftarrow\operatorname{argmin}_{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}\,f(\Pi_{\mathcal{S}_{1:d}})$ $\triangleright$ flag trick (11) + optimization (Alg. 1)
@@
  $g_{k}^{*}\leftarrow\operatorname{fit}(\Pi_{\mathcal{S}_{k}^{*}}X,Y)$ $\triangleright$ learning on $q_{k}$-dimensional projected data
  $Y_{k}^{*}\leftarrow g_{k}^{*}(\Pi_{\mathcal{S}_{k}^{*}}X)$ $\triangleright$ prediction on $q_{k}$-dimensional projected data
@@
$Y^{*}\leftarrow\operatorname{ensembling}({Y}_{1}^{*},\dots,{Y}_{d}^{*})$ $\triangleright$ weighted predictions
@@
Figure 2: Illustration of the flag trick methodology. 1) We start with a subspace learning problem: $\operatorname*{arg\,min}_{\mathcal{S}\in\operatorname{Gr}(p,q)}f(\Pi_{\mathcal{S}})$. Trying different $q$’s yields in general non-nested subspaces, which raises an issue of consistency between data representations. 2) We convert the subspace learning problem into a nested subspace learning problem via the flag trick: $\operatorname*{arg\,min}_{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}f(\frac{1}{d}\sum_{k=1}^{d}\Pi_{\mathcal{S}_{k}})$. We run a steepest descent on flag manifolds (Algorithm 1) and get a flag of nested subspaces. 3) We fit a machine learning algorithm (regression, classification, etc.) to the projected data at each dimension $q_{k}\in q_{1:d}$. We aggregate the estimators via ensembling methods (hard voting, soft voting, etc.) and get improved predictions.
Algorithm 2 is a general proposition of multilevel machine learning with flags, but many other uses of the optimal flag $\mathcal{S}_{1:d}^{*}$ are possible, depending on the application. For instance, one may directly use the reweighted data matrix $\Pi_{\mathcal{S}_{1:d}^{*}}X$ as an input to the machine learning algorithm. This enables to fit only one model instead of $d$. One can also simply analyze the projected data qualitatively via scatter plots or reconstruction plots as evoked in Section 1. The nestedness will automatically bring consistency contrarily to non-nested subspace methods, and therefore improve interpretability. Finally, many other ideas can be borrowed from the literature on subspace clustering and flag manifolds (draper_flag_2014; launay_mechanical_2021; ma_flag_2021; mankovich_flag_2022; mankovich_chordal_2023; mankovich_fun_2024; mankovich_flag_2025), for instance the computation of distances between flags coming from different datasets as a multilevel measure of similarity between datasets.
Remark 12 (Choice of Signature)
In the following experiments, the flag signatures are chosen heuristically. Although important, we believe that such a choice must be thoroughly addressed for each subspace learning problem, which leads us to leave that for future research. Alternative methodologies for the automatic choice of dimensions are discussed in Section 5. One could notably investigate penalties on the spectral gaps in the same spirit as szwagier_eigengap_2025—where the penalty factor would be tuned according to the downstream machine learning task. Beyond spectral gaps, the signature could be selected by starting from a flag of signature $(p,(1,2,\dots,p-1))$ and removing the nested subspaces that do not significantly increase the accuracy on the downstream machine learning task. For the specific example of subspace averaging, a hyperparameter-free signature selection method extending the order fitting rule of santamaria_order_2016 could be investigated via the flag trick.
@@ 4.1 Outline and Experimental Setting
The general methodology to compare Grassmann-based methods to flag-based methods is the following one. For each experiment, we first choose a flag signature $\penalty 10000\ {q_{1:d}\coloneqq(q_{1},\dots,q_{d})}$, then we run independent optimization algorithms on $\operatorname{Gr}(p,q_{1}),\dots,\operatorname{Gr}(p,q_{d})$ (10) and finally we compare the optimal subspaces $\mathcal{S}_{k}^{*}\in\operatorname{Gr}(p,q_{k})$ to the optimal flag of subspaces $\mathcal{S}_{1:d}^{*}\in\operatorname{Fl}(p,q_{1:d})$ obtained via the flag trick (11). To show the nestedness issue in Grassmann-based methods, we compute the subspace distances $\Theta(\mathcal{S}_{k}^{*},\mathcal{S}_{k+1}^{*})_{k=1\dots d-1}$, where $\Theta$ is the generalized Grassmann distance of ye_schubert_2016. It consists in the $\ell^{2}$-norm of the principal angles, which can be obtained from the singular value decomposition (SVD) of the inner-product matrices ${U_{k}}^{\top}{U_{k+1}}$$U_{k}\in\operatorname{St}(p,q_{k})$ being an orthonormal basis of $\mathcal{S}_{k}^{*}$.
Regarding the implementation of the steepest descent algorithm on flag manifolds (Algorithm 1), we develop a new class of manifolds in PyManOpt (boumal_manopt_2014; townsend_pymanopt_2016), and run their SteepestDescent algorithm. Our implementation of the Flag class is based on the Stiefel representation of flag manifolds, detailed in Section 2, with the retraction being the polar retraction. For the computation of the gradient, we use automatic differentiation with the autograd package. We could derive the gradients by hand from the expressions we get, but we use automatic differentiation as strongly suggested in PyManOpt’s documentation. Finally, the real datasets and the machine learning methods used in the experiments can be found in scikit-learn (pedregosa_scikit-learn_2011).
@@ 4.2 The Flag Trick for Robust Subspace Recovery Problems
As shown in introduction and particularly in Figure 1, robust subspace recovery (RSR) methods based on Grassmannian optimization might suffer from the nestedness issue, which notably yields inconsistent data representations across different dimensions (as seen on the left part of the figure, where outliers somewhat swap position with inliers from 1D to 2D). The RSR overview by lerman_overview_2018 particularly emphasizes the need for nested subspace methods, although it leaves this as an open perspective. In this subsection, we address the nestedness issue in robust subspace recovery by applying the flag trick (Definition 10) to the problem of absolute deviation minimization (maunu_well-tempered_2019).
Robust subspace recovery is an outlier-robust extension of classical dimension reduction methods such as PCA. Let us consider a dataset that is a union of inliers and outliers—the inliers are assumed to lie near a low-dimensional subspace $\mathcal{S}$ while the outliers live in the ambient space. The aim of RSR is to recover $\mathcal{S}$. Without further specifications, the RSR problem might not be well-posed. Therefore, the works in this domain often have to make some assumptions on the inlier and outlier distributions in order to obtain convergence and recovery guarantees. For instance, in lerman_overview_2018, it is assumed that the inliers “fill” the lower-dimensional subspace and that the outliers are not much “aligned”; this is rigorously defined in lerman_robust_2015 and maunu_well-tempered_2019 through permeance and alignment statistics.
A generative model—the haystack model—following these assumptions is introduced in lerman_robust_2015. It assumes an isotropic Gaussian distribution on the subspace for the inliers and an isotropic Gaussian distribution on the (full) ambient space for the outliers. A more realistic model—the generalized haystack model—is introduced in maunu_well-tempered_2019 to circumvent the simplistic nature of the haystack model. This one assumes general (anisotropic) Gaussian distributions for the inliers and outliers. [^3] The generalized haystack model is more-precisely defined as follows: the $n_{\mathrm{in}}$ inliers are i.i.d sampled from a sub-Gaussian distribution $\mathcal{N}\left(0,\Sigma_{\mathrm{in}}/q\right)$, with $\Sigma_{\mathrm{in}}=U\Lambda U^{\top}$, $U\in\operatorname{St}(p,q)$ and $\Lambda\in\operatorname{diag}\left(\mathbb{R}^{q}\right)$, while the $n_{\mathrm{out}}$ outliers are i.i.d sampled from a sub-Gaussian distribution $\mathcal{N}\left(0,\Sigma_{\mathrm{out}}/\operatorname{rk}(\Sigma_{\mathrm{out}})\right)$, with $\Sigma_{\mathrm{out}}\in\operatorname{Sym}(p)$. The (original) haystack model is a particular case of the generalized haystack model with isotropic outliers and (subspace-)isotropic inliers, i.e., $\Sigma_{\mathrm{out}}=\sigma^{2}_{\mathrm{out}}\,I_{p}/p$ and $\Sigma_{\mathrm{in}}=\sigma^{2}_{\mathrm{in}}\,UU^{\top}/q$. This makes the learning harder, since the anisotropy may keep the inliers from properly permeating the low-dimensional subspace—as discussed in Section 1. Therefore, one has to make some stronger assumptions on the inlier-outlier ratio and the covariance eigenvalues distributions to derive some convergence and recovery guarantees.
@@
The inlier distribution in the haystack model follows the isotropic PPCA model (bouveyron_hddc_2007; bouveyron_intrinsic_2011), while it follows the PPCA model (tipping_probabilistic_1999) in the case of the generalized haystack model. Both models are a special case of the principal subspace analysis models (szwagier_curse_2025). However, as argued in szwagier_curse_2025, while the haystack model is parameterized with Grassmannians, the generalized haystack model—which has more degrees of freedom accounting for the anisotropy—is parameterized with Stiefel manifolds. Therefore, from a statistical modeling perspective, it only makes sense to conduct subspace learning experiments on the haystack model and not the generalized one.